# Explain with Visual Keypoints Like a Real Mentor! A Benchmark for Multimodal Solution Explanation

### v3

@@ Abstract
With the rapid advancement of mathematical reasoning capabilities in Large Language Models (LLMs), AI systems are increasingly being adopted in educational settings to support students’ comprehension of problem-solving processes. However, a critical component remains underexplored in current LLM-generated explanations: multimodal explanation. In real-world instructional contexts, human tutors routinely employ visual aids, such as diagrams, markings, and highlights, to enhance conceptual clarity. To bridge this gap, we introduce the multimodal solution explanation task, designed to evaluate whether models can identify visual keypoints, such as auxiliary lines, points, angles, and generate explanations that incorporate these key elements essential for understanding. To evaluate model performance on this task, we propose ME2, a multimodal benchmark consisting of 1,000 math problems annotated with visual keypoints and corresponding explanatory text that references those elements. Our empirical results show that, aside from recent large-scale open-source and closed-source models, most generalist open-source models, and even math-specialist models, struggle with the multimodal solution explanation task. This highlights a significant gap in current LLMs’ ability to reason and explain with visual grounding in educational contexts. We expect that the multimodal solution explanation task and the ME2 dataset will catalyze further research on LLMs in education and promote their use as effective, explanation-oriented AI tutors. Code and data will be released publicly.
@@ 1 Introduction
The traditional one-to-many educational model (i.e., one teacher for multiple students) is gradually transitioning to one-to-one personalized tutoring systems and online learning (Mukul and Büyüközkan 2023). Recent developments in Multimodal Large Language Models (MLLMs) have opened new opportunities for effective learning, such as estimating question difficulty (Park et al. 2024), assisting teachers in curriculum planning (Hu et al. 2024), and supporting interactive tutoring systems (Chevalier et al. 2024). In particular, numerous studies (Liu et al. 2023; Uesato et al. 2022; Lu et al. 2023) have focused on enhancing the mathematical reasoning abilities of MLLMs. As a result, MLLMs have led many students to use them as tools when faced with mathematical questions (Pardos and Bhandari 2024).
@@
However, from a student’s perspective, relying solely on the reasoning footprints of MLLMs may not always be the best way to understand problems (Pardos and Bhandari 2024; Jia et al. 2024). One might wonder what distinguishes a broadly comprehensible explanation from a solution that merely yields the correct answer, for either a human or a model? A critical factor is the use of visual cues.
@@
In actual educational settings, Dual Coding Theory (DCT) naturally occurs, providing effective learning opportunities for students (Paivio 2013, 1990). According to DCT, combining verbal and visual information enhances student comprehension (Clark and Paivio 1991). As illustrated in Figure 1, human mentors often use visual scaffolding, such as annotated diagrams or highlighted keypoints on a blackboard, to foster intuitive understanding (Arcavi 2003; Stylianou 2010; Lee, Park, and Park 2024). In contrast, current AI models lack the capacity to generate such visual explanations. Moreover, existing datasets for training and evaluation are insufficient for developing models capable of producing these forms of multimodal instructional support (Hendrycks et al. 2021; Lu et al. 2023; Wang et al. 2024a).
@@
To address these limitations, we introduce multimodal solution explanation, a novel task designed to evaluate and advance models’ ability to generate educationally effective, visually grounded mathematical explanations. To benchmark performance on multimodal solution explanation, we propose Multimodal Explanations for Mathematics Education (ME2) benchmark. The multimodal solution explanation task requires (1) identifying visual keypoints that are not present in the original problem but crucial for understanding (e.g., lines, angles, annotations) and (2) generating explanatory text that explicitly refers to them. The ME2 includes not only the original problem images and text, but also annotations of the visual keypoints that illustrate how the solution images differ from the originals, as well as keypoint-based solution explanations. Notably, we emphasize the educational value of ME2, as it is derived from real-world instructional contexts.
Experiments on ME2 demonstrate that, while closed-source models show potential on multimodal solution explanation, most current open-source generalist models still struggle to reliably identify visual keypoints. Even mathematical models, which are fine-tuned on mathematical domains, remain primarily focused on problem-solving and often do not produce sufficiently structured outputs to highlight critical visual details. These findings underscore a critical opportunity to steer future model development toward more effective visual grounding in educational contexts. We believe that ME2 will catalyze research aimed at strengthening visual grounding for mathematical reasoning and advancing multimodal solution explanation models capable of serving as effective, student-friendly educational mentors.
@@ 2 ME2 Benchmark
ME2 is a multimodal solution explanation benchmark consisting of 1,000 instances. Each of which contains a problem text ($T_{p}$), a problem image ($I_{p}$), an explanatory solution text ($T_{s}$), a solution image ($I_{s}$), and visual keypoints ($VK$) that highlight how the solution image differs from the original, along with a concise summary of the explanation ($T_{s}^{tldr}$) to fix the solution direction and assess alignment between keypoints and their corresponding explanations, even when multiple valid solution exist.
To create a benchmark that can assess the multimodal solution explanation capabilities of MLLMs, we define the visual keypoint $VK$ and summary of the explanation $T_{s}^{tldr}$ in Section 2.1. An overview of ME2 and its construction is illustrated in Figure 2.
@@ 2.1 Benchmark Construction
Table 1: Statistics of the ME2 benchmark, including problem subjects and types, visual keypoints, and word counts.
Annotation Process.
To create the textual visual keypoints, we streamline the simple yet labor-intensive task of comparing problem and solution images by using GPT-4o (Achiam et al. 2023) as an auxiliary tool. The model produces an initial set of keypoints $\{vk^{ai}_{1},\dots,vk^{ai}_{n}\}\in VK^{ai}$, which four annotators, each holding a bachelor’s degree in science or engineering, verify and refine for accuracy and compliance with our formatting guidelines:
Once the verified keypoints are fixed, we generate a brief, keypoint-aligned summary of each solution text, $T_{s}^{tldr}$. This summary captures the essential reasoning trajectory so that, even when a problem permits multiple valid approaches, a single, unambiguous set of visual keypoints remains authoritative. As with the keypoints, AI tool produces the initial draft, and human annotators review and refine it for clarity and consistency. Consequently, each ME2 instance is represented as $\langle T_{p},I_{p},T_{s},VK,T_{s}^{\text{tldr}}\rangle$. Further details are provided in the Supplementary Materials.
@@
2.2 Data Analysis
The ME2 benchmark consists of 1,000 problem–solution pairs: 763 (76.3%) geometry problems and 237 (23.7%) graph problems. Among these, 605 (60.5%) are multiple-choice questions and 395 (39.5) are short-answer questions. It spans 17 chapters (see Figure 4) and 33 sections (see Supplementary Materials). On average, each sample contains about 3.8 visual keypoints $VK$, derived from annotations. These keypoints fall into four main categories: points, lines, regions, and symbols. The symbol category is further divided into parallel marks, equal-length marks, right-angle marks, and length-label marks. Additional statistical details about visual keypoints $VK$, and length statistics for the problem text $T_{p}$, the solution text $T_{s}$, and the visual keypoint components $vk_{n}$ are provided in Table 1.
3 Task Definition
We propose two tasks to evaluate a model’s multimodal solution explanation capability. Multimodal solution explanation requires identifying visual keypoints useful for subsequent explanation and generating explanations that explicitly reference them. These tasks offer a chance to assess how well the model understands multimodal educational elements and produces informative explanations along with visual keypoints. The two tasks are illustrated in Figure 3.
Visual Keypoint Identification.
The first task evaluates the model’s ability to identify visual keypoints that are crucial for comprehension. To prevent ambiguity arising from multiple solution paths to each problem, we supply the model with a summary of solution text $T_{s}^{tldr}$, ensuring it focuses on a single solution path and enabling precise evaluation of its ability to identify the corresponding visual keypoints. Given a problem image $I_{p}$, its text $T_{p}$, the correct answer, and the solution summary $T_{s}^{tldr}$, the model must select, from five candidate sets, the visual keypoints ($VK$) essential for understanding. The four distractor sets are constructed as follows: (1) $VK$ from a problem whose text is semantically similar to $T_{p}$; (2) $VK$ from a problem whose solution summary resembles $T_{s}$; (3) $VK$ from a problem whose own keypoints closely match the target $VK$ (4) $VK$ from a randomly selected problem. We compute text similarity using Qwen3 Embedding (Zhang et al. 2025).
Keypoint-based Explanation Generation.
The second task evaluates whether the model can effectively generate explanatory text grounded in the appropriate visual keypoints. We provide the set of visual keypoints ($VK$) to eliminate ambiguity and guide the model toward a single correct reasoning path. Given a problem consisting of an image $I_{p}$, text $T_{p}$, and problem’s answer, along with the visual keypoints $VK$, the model is required to produce a solution explanation $T_{s}$ that refers to the relevant visual elements.
@@ 3 Task Definition
Table 2: Experimental results on the Solution Recognition toy task from ME2. The best scores are shown in bold, and the second-best scores are underlined.
4 Experiments
Models.
We evaluate three categories of MLLMs: Open-source Generalist (OG), Open-source Mathematical specialist (OM), and Closed-source Generalist (CG). The OG models includes widely used models such as Molmo 7B (Deitke et al. 2024), LLaVA-1.6 7B (Liu et al. 2024), Qwen 2-VL (Wang et al. 2024b), and Qwen 2.5-VL in both 7B and 72B versions (Bai et al. 2025). The CG models comprise GPT-4o (Achiam et al. 2023) and Gemini 2.0 Flash (Google DeepMind 2024). Finally, the OM models, specialized for mathematics, include Math-PUMA 7B (Zhuang et al. 2024), URSA 8B (Luo et al. 2025), and Math-LLaVA 13B (Shi et al. 2024). Details of the experimental setup and prompts are provided in the Supplementary Materials.
4.1 Toy: Solution Recognition
Inspired by real-world tutoring settings where tutors typically know the solution path, the multimodal solution explanation tasks are designed under the assumption that the model has access to the correct solution, either in the form of a summary solution or a set of visual keypoints. To assess whether the model understands how to solve problems, we first conduct a preliminary study on ME2 before proceeding with the multimodal solution explanation task.
Metrics.
ME2 consists of both multiple-choice and short-answer problems. We report accuracy following the MathVista evaluation protocol (Lu et al. 2023).
Results.
Table 2 shows the accuracy of each model on the Solution Recognition toy task. The 7B OG baseline struggles, while the 72B model performs second best. Somewhat unexpectedly, the OM models perform worse than the OG models, likely due to aggressive fine-tuning that hinders instruction-following capabilities. Math-PUMA, despite being fine-tuned for mathematics on Qwen2-VL, also underperforms compared to the original models. Among the CG baselines, GPT-4o struggled similarly to open-source models, whereas Gemini achieved the best performance overall. These results indicate that most open-source models struggle to recognize the correct solution on ME2, even before performing the multimodal solution explanation task
@@ 4.2 Visual Keypoint Identification
We evaluate baselines’ VK identification performance using accuracy under a multiple-choice setting.
@@ Results.
Table 3 summarizes performance on the VK Identification task. Among the OG baselines, the 7B models still face considerable difficulty, while the 72B model remains the second-best performer, similar to the Solution Recognition toy task. In contrast, the OM models perform at or below chance level (Acc = 0.20), even worse than on the toy task, suggesting a fundamental inability to detect relevant visual cues or a loss of instruction-following capability. Among the CG models, GPT-4o shows the most noticeable improvement despite its earlier underperformance on the toy task, while Gemini consistently achieves the highest performance. These findings indicate that although most open-source models still struggle despite having access to the solution, CG models and larger OG models can reasonably identify the visual keypoints.
@@ 4.3 Keypoint-based Explanation Generation
We evaluate the quality of the explanation using three criteria: (1) Correctness – whether the model’s reasoning is logically sound and leads to a valid solution; (2) Fidelity – whether the explanation aligns with the reasoning and intent of the reference, regardless of surface form; (3) Referencing – whether the explanation refers to the same key visual components (e.g. points, lines, etc) as the reference. Each criterion is rated on a 5-point Likert scale. We report results from both human evaluators (Zheng et al. 2023) and an LLM-based evaluator (Achiam et al. 2023). In addition, we report text similarity metrics, including BLEU, ROUGE, METEOR, and BERTScore (Papineni et al. 2002; Lin 2004; Banerjee and Lavie 2005; Zhang et al. 2019), with further details provided in the Supplementary Materials.
@@
[…]

### v4

@@ Abstract
With the rapid advancement of mathematical reasoning capabilities in Large Language Models (LLMs), AI systems are increasingly being adopted in educational settings to support students’ comprehension of problem-solving processes. However, a critical component remains underexplored in current LLM-generated explanations: multimodal explanation. In real-world instructional contexts, human tutors routinely employ visual aids, such as diagrams, markings, and highlights, to enhance conceptual clarity. To bridge this gap, we introduce the multimodal solution explanation task, designed to evaluate whether models can identify visual keypoints, such as auxiliary lines, points, angles, and generate explanations that incorporate these key elements essential for understanding. To evaluate model performance on this task, we propose ME2, a multimodal benchmark consisting of 1,000 math problems annotated with visual keypoints and corresponding explanatory text that references those elements. Our empirical results show that current models struggle to identify visual keypoints. In the task of generating keypoint-based explanations, open-source models also face notable difficulties. This highlights a significant gap in current LLMs’ ability to perform mathematical visual grounding, engage in visually grounded reasoning, and provide explanations in educational contexts. We expect that the multimodal solution explanation task and the ME2 dataset will catalyze further research on LLMs in education and promote their use as effective, explanation-oriented AI tutors.
Archivehttps://me2-benchmark.github.io
@@ 1 Introduction
The traditional one-to-many educational model (i.e., one teacher for multiple students) is gradually transitioning to one-to-one personalized tutoring systems and online learning (mukul2023digital). Recent developments in Multimodal Large Language Models (MLLMs) have opened new opportunities for effective learning, such as estimating question difficulty (park2024large), assisting teachers in curriculum planning (hu2024teaching), and supporting interactive tutoring systems (chevalier2024language). In particular, numerous studies (liu2023mathematical; uesato2022solving; lu2023mathvista) have focused on enhancing the mathematical reasoning abilities of MLLMs. As a result, MLLMs have led many students to use them as tools when faced with mathematical questions (pardos2024chatgpt).
@@
However, from a student’s perspective, relying solely on the reasoning footprints of MLLMs may not always be the best way to understand problems (pardos2024chatgpt; jia2024comparison). One might wonder what distinguishes a broadly comprehensible explanation from a solution that merely yields the correct answer, for either a human or a model? A critical factor is the use of visual cues.
@@
In actual educational settings, Dual Coding Theory (DCT) naturally occurs, providing effective learning opportunities for students (paivio2013imagery; paivio1990mental). According to DCT, combining verbal and visual information enhances student comprehension (clark1991dual). As illustrated in Figure 1, human mentors often use visual scaffolding, such as annotated diagrams or highlighted keypoints on a blackboard, to foster intuitive understanding (arcavi2003role; stylianou2010teachers; lee2024vistavisualintegratedtailored). In contrast, current AI models lack the capacity to generate such visual explanations. Moreover, existing datasets (hendrycks2021measuring; lu2023mathvista; wang2024measuring) focus solely on problem-solving and overlook educational objectives, making them insufficient for developing models capable of providing such forms of multimodal instructional support.
@@
To address these limitations, we introduce multimodal solution explanation, a novel task that aims to enhance models’ capacity to generate educationally effective and visually grounded mathematical explanations. In this task, models are required to (1) identify visual keypoints that are not present in the original problem but are crucial for understanding (e.g., lines, angles, annotations), and (2) generate explanatory text that explicitly refers to them. To benchmark performance on multimodal solution explanation, we propose Multimodal Explanations for Mathematics Education (ME2) benchmark. The ME2 includes not only the problem and its solution, but also annotations of the visual keypoints that serve as visual cues necessary to explain the solution, as well as keypoint-based explanatory texts aligned with them. Notably, we emphasize that ME2 goes beyond simple problem-solving, offering educational value in addressing the previously unexplored dimension of multimodal solution explanation.
Experiments on ME2 demonstrate that current MLLMs struggle to reliably identify visual keypoints. While closed-source models show potential in generating explanations grounded in visual keypoints, open-source generalist and math-specialized models show limited ability in this aspect. This suggests that current models largely fail to achieve robust mathematical visual grounding and visually grounded reasoning in educational contexts. We believe that ME2 will catalyze research toward strengthening mathematical visual grounding and reasoning, and advancing models that can serve as effective and student-friendly educational mentors.
@@
2 Related Works
Language Models for Education.
Recent advances in Large Language Models (LLMs) (NEURIPS2020_1457c0d6) have sparked significant interest in educational applications, particularly in personalized problem recommendation (park2024large), automated tutoring (chevalier2024language), and the provision of tailored feedback and customized curricula (hu2024teaching; macina2023mathdial; feng2023citing). For effective education, research suggests that combining textual and visual information enhances comprehension and memory more effectively than using text alone (arcavi2003role; stylianou2010teachers; lee2024vistavisualintegratedtailored). As clark1991dual explains, dual representations supply multiple retrieval cues and cultivate richer mental models. Building on this insight, we introduce the multimodal solution explanation task to enable LLMs to offer learners more comprehensive learning opportunities. This task enables the model to pinpoint the visual keypoints essential for students’ comprehension and to generate explanations grounded in those keypoints, thereby delivering more comprehensive educational support and ultimately improving the overall quality of their learning experience.
Mathematical Benchmarks.
Current LLMs show strong performance on mathematical problems, making them valuable tools for students (zhuang2024math; luo2025ursa). To evaluate these models, traditional mathematical benchmarks (cobbe2021training; hendrycks2021measuring) have been crucial in assessing reasoning capabilities. With the growing multimodal capabilities of LLMs, benchmarks such as MathVista (lu2023mathvista), Math-Vision (wang2024measuring), and MathVerse (zhang2024mathverse) have extended this evaluation to image-based math problems. Recent efforts like OlympiadBench (he2024olympiadbench) and MM-MATH (sun2024mm) further assess models not just on final answers but also on their reasoning processes. However, most existing benchmarks focus solely on problem-solving, overlooking educational objectives. To address this gap, we introduce ME2, which advances beyond problem-solving to evaluate a model’s capacity to generate visually and logically coherent explanations and key visual cues that support effective instructional use.
@@ 3 ME2 Benchmark
ME2 is a multimodal solution explanation benchmark consisting of 1,000 instances. Each of which contains a problem text ($T_{p}$), a problem image ($I_{p}$), an explanatory solution text ($T_{s}$), a solution image ($I_{s}$), and visual keypoints ($VK$) that newly highlight elements crucial for understanding (e.g., lines, angles), and a concise explanation summary ($T_{s}^{tldr}$) to anchor the model’s explanatory solution direction. To create a benchmark that can assess the multimodal solution explanation capabilities of MLLMs, we define the visual keypoint $VK$ and summary of the explanation $T_{s}^{tldr}$ in Section 3.1. An overview of ME2 and its construction is illustrated in Figure 2.
@@ 3.1 Benchmark Construction
Annotation Process.
To create the textual-form visual keypoints, we streamline the simple yet labor-intensive task of comparing problem and solution images by using GPT-4o (achiam2023gpt) as an auxiliary tool. The model produces an initial set of keypoints $\{vk^{\text{ai}}_{1},\dots,vk^{\text{ai}}_{n}\}\in VK^{\text{ai}}$, which four annotators, each holding a bachelor’s degree in science or engineering, verify and refine for precision and consistency with our annotation guidelines:
Once the verified keypoints are fixed, human annotators and the AI tool jointly generate a brief, keypoint-aligned summary of each solution text ($T_{s}^{tldr}$). Since explanations may follow multiple valid paths (see Appendix), this summary anchors a single solution direction during model explanation generation, ensuring an unambiguous consensus set of visual keypoints. Finally, the entire benchmark was translated from Korean to English using an AI tool and then reviewed by two bilingual annotators. From a 10% subset, annotators achieved substantial agreement, with Cohen’s $\kappa$ of 0.84 (cohen1960coefficient), indicating strong reliability. Consequently, each ME2 instance is represented as $\langle T_{p},I_{p},T_{s},VK,T_{s}^{\text{tldr}}\rangle$.
@@
Table 1: Statistics of the ME2 benchmark, including problem subjects, types, and instance word counts.
3.2 Data Analysis
The ME2 benchmark consists of 1,000 problem–solution pairs: 763 (76.3%) geometry problems and 237 (23.7%) graph problems. Among these, 605 (60.5%) are multiple-choice questions and 395 (39.5%) are short-answer questions. It spans 17 chapters (see Figure 4) and 33 sections (see Appendix). On average, each sample contains about 3.8 visual keypoints $VK$, derived from annotations. These keypoints fall into four main categories: points, lines, regions, and symbols. The symbol category is further divided into parallel marks, equal-length marks, right-angle marks, and length-label marks. Additional statistical details about visual keypoints $VK$, and length statistics for the problem text $T_{p}$, the solution text $T_{s}$, and the visual keypoint components $vk_{n}$ are provided in Table 1.
4 Task Definition
We propose two tasks to evaluate a model’s multimodal solution explanation capability, as illustrated in Figure 3. The first task requires (1) identifying visual keypoints useful for subsequent explanation and (2) generating explanations that explicitly reference them. For robust evaluation, we structured the tasks to isolate perceptual and reasoning subskills. Although the design abstracts away some real-world complexity, this two-stage setup still provides a clear and measurable step toward unified, open-ended reasoning.
Visual Keypoint Identification.
The first task evaluates the model’s ability to identify visual keypoints that are crucial for comprehension. Since a problem may have multiple valid solutions and the corresponding visual keypoints can vary, we provide the model with a solution summary ($T_{s}^{tldr}$) that anchors a single explanatory direction. To ensure that the evaluation focuses on keypoint identification rather than problem-solving, the correct answer is also provided. Additionally, because current models cannot reliably generate valid keypoints and open-ended scoring is ambiguous, we adopt a multiple-choice format for robust evaluation.
Given a problem image $I_{p}$, its text $T_{p}$, the correct answer, and the solution summary $T_{s}^{tldr}$, the model must select, from five candidate sets, the visual keypoints ($VK$) essential for understanding. The four distractor sets are constructed as follows: (1) $VK$ from a problem whose text is semantically similar to $T_{p}$; (2) $VK$ from a problem whose solution summary resembles $T_{s}$; (3) $VK$ from a problem whose own keypoints closely match the target $VK$ (4) $VK$ from a randomly selected problem. We compute text similarity using Qwen3 Embedding (qwen3embedding).
Keypoint-based Explanation Generation.
The second task evaluates whether the model can effectively generate explanatory text grounded in the appropriate visual keypoints. As in the first task, we provide visual keypoints ($VK$) to guide the model toward a single reasoning path and the correct answer to focus evaluation on keypoint-aligned explanation generation rather than problem-solving.
Given a problem consisting of an image $I_{p}$, text $T_{p}$, and problem’s answer, along with the visual keypoints $VK$, the model is required to produce a solution explanation $T_{s}$ that refers to the relevant visual elements.
@@ 4 Task Definition
5 Experiments
Models.
We evaluate three categories of MLLMs: (1) generalist models, including Molmo 7B (deitke2024molmo), LLaVA-1.6 7B (liu2024improved), Qwen2-VL 7B (wang2024qwen2), and Qwen2.5-VL 7B & 72B (bai2025qwen2); (2) math-specialized models, including Math-PUMA 7B (zhuang2024math), URSA 8B (luo2025ursa), and Math-LLaVA 13B (shi-etal-2024-math); and (3) proprietary models, GPT-4o (achiam2023gpt) and Gemini 2.0 Flash (google2024gemini2). Details of the experimental setup and prompts are provided in the Appendix.
5.1 Toy: Solution Recognition
The multimodal solution explanation tasks are designed to evaluate specific abilities rather than general problem-solving skills. To examine whether the model genuinely understands how to solve problems, we first perform a preliminary study on ME2 prior to the multimodal solution explanation task.
Metrics.
ME2 consists of both multiple-choice and short-answer problems. We report accuracy following the MathVista evaluation protocol (lu2023mathvista).
Results.
Table 2 shows the accuracy of problem-solving. The 7B generalist baseline struggles, while the 72B model performs second best. Somewhat unexpectedly, the math-specialized models perform worse than the generalist models, likely due to hindered instruction-following capabilities. Among the proprietary baselines, GPT-4o struggled similarly to open-source models, whereas Gemini achieved the best performance overall. These results indicate that most MLLMs struggle to recognize the correct solution on ME2, even before performing the multimodal solution explanation task.
@@ Results.
Table 2: Experimental results on the Solution Recognition toy task from ME2. Models are grouped into three categories: generalist models (top), math-specialized models (middle), and proprietary models (bottom). The best scores are in bold, and the second-best scores are underlined.
@@ 5.2 Visual Keypoint Identification
We evaluate baseline performance using accuracy in a multiple-choice setting.
@@ Results.
Table 3 summarizes performance on the visual keypoint identification task. The 7B generalist models struggle, while the 72B model remains the second-best performer. In contrast, math-specialized models perform near chance level (Acc = 0.20), indicating severe difficulty in identifying visual cues. Among proprietary models, Gemini achieves the highest performance. Overall, most models struggle to identify visual keypoints even with access to the solution, though proprietary ones perform relatively better.
Since visual keypoint identification was evaluated under the assumption that models can already perform problem-solving, Table 4 reports success rates for each task and for both together. Only 23%, 10%, and 4% of proprietary, generalist, and math-specialized models succeed on both. This result highlights that real educational use remains challenging and that improving problem-solving ability is essential alongside visual keypoint identification.
@@
Model PS Only VKI Only PS $\cap$ VKI
$\text{Qwen2.5-VL}^{\text{7B}}$ 18.9% 29.8% 10.5%
Math-PUMA 19.7% 15.5% 4.6%
Gemini 2.0 F 20.5% 34.5% 23.1%
@@ 5.3 Keypoint-based Explanation Generation
We evaluate the quality of the explanation using three criteria: (1) Correctness – whether the model’s reasoning is logically sound and leads to a valid solution; (2) Fidelity – whether the explanation aligns with the reasoning and intent of the reference, regardless of surface form; (3) Referencing – whether the explanation refers to the same key visual components (e.g. points, lines, etc) as the reference. Each criterion is rated on a 5-point Likert scale. We report results from both human evaluators (zheng2023judging) and an LLM-based evaluator using GPT-4o (achiam2023gpt). In addition, we report text similarity metrics, including BLEU, ROUGE, METEOR, and BERTScore (papineni2002bleu; lin2004rouge; banerjee2005meteor; zhang2019bertscore), with further details provided in the Appendix.
@@
$\text{Qwen2.5-VL}^{\text{7B}}$ 3.005 2.375 3.132
$\text{Qwen2.5-VL}^{\text{72B}}$ 3.397 3.048 3.533
@@
Results.
Table 5 presents the results of the explanation generation task. While most models achieve reasonable Correctness, many fail to follow the intended reasoning path (Fidelity) or reference the given keypoints (Referencing). Generalist models struggle overall, though the Qwen2.5-VL series shows size-dependent improvement. Math-specialized models still fail to produce coherent or instruction-following explanations. In contrast, proprietary models achieve the highest scores, demonstrating stronger abilities in generating well-grounded explanations.
As shown in Table 6, human evaluation exhibits strong correlations with LLM judgments, as indicated by the Spearman coefficients (zar2005spearman) (0.770 for Correctness, 0.783 for Fidelity, 0.788 for Referencing; all p<<0.05). These results show that although most open-source models still struggle, proprietary models and more recent generalist models can generate appropriate explanations.
6 Analyses
6.1 Qualitative Analysis
To analyze how the three categories of models differ in their outputs, we examined the results from representative models in each category: Qwen2.5-VL 7B (generalist), Math-PUMA (specialist), and Gemini 2.0 Flash (proprietary).